OpenStudy (anonymous):

\[\int_{9}^{16} \frac{\sqrt(x)}{x-4} dx\] \[u=\sqrt(x)\] \[du=\frac 1 2 \frac{1}{\sqrt(x)}dx\] \[2\sqrt(x) du=dx\] \[2udu=dx\] \[\int_{3}^{4} \frac{u^2}{u^2-4} du\] \[2\int_{3}^{4} 1+\frac{4}{(u-2)(u+2)} du\] \[2\int_{3}^{4} du +2\int_{3}^{4}\frac{4}{(u-2)(u+2)} du\] \[2\int_{3}^{4} du +8\int_{3}^{4} \frac{1}{(u-2)} - \frac{1}{(u+2)}du\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Are we on the right track?

OpenStudy (anonymous):

yes, there is a slight typo in the third line where the two disappears, but it comes back, so you are fine

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[2\int_{3}^{4} du +8\int_{3}^{4} \frac{1}{(u-2)}du - 8\int_{3}^{4} \frac{1}{(u+2)}du\]

OpenStudy (anonymous):

let me check the constants carefully

OpenStudy (anonymous):

ok there is a mistake in the constants

OpenStudy (anonymous):

I knew it!

OpenStudy (anonymous):

\[\frac{4}{(x+2)(x-2)}=\frac{1}{x-2}-\frac{1}{x+2}\] you used the 4 twice it looks like

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

how so as in why it is true that \[\frac{4}{(x-2)(x+2)}=\frac{1}{x-2}-\frac{1}{x+2}\] or how so as in how did you use the 4 twice?

OpenStudy (anonymous):

how did i use the 4 twice

OpenStudy (anonymous):

\[\frac{4}{(x-2)(x+2)}=\frac{1}{x-2}-\frac{1}{x+2}\] is the above statement correct though?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now the 4 is gone right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and therefore you do not get an 8 out front, but rather the 2 you started with that is the mistake

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks....that 4 looked suspicious

OpenStudy (anonymous):

\[2\int_{3}^{4} du +2\int_{3}^{4}\frac{4}{(u-2)(u+2)} du\] \[=2\int_{3}^{4} du +2\int_{3}^{4} \frac{1}{(u-2)} - \frac{1}{(u+2)}du\]

OpenStudy (anonymous):

thanks for catching that!

OpenStudy (anonymous):

been at this all day?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

no ...i did some other problems in the mean time...but yeah...I kept making small mistakes and confusing myself

OpenStudy (anonymous):

that's when I decided to make the face of "confusion" my profile pic

OpenStudy (anonymous):

a lot of book keeping easy to make a mistake

OpenStudy (anonymous):

lol